What is the standard form of the conic with the equation:
3x^2+5y^2-12x+30y+42=0

Answer by stanbon(75887)

What is the standard form of the conic with the equation:
3x^2+5y^2-12x+30y+42=0
---
Complete the square on the x-terms and on the y-terms separately:
3x^2 - 12 x + ? + 5y^2 + 30y + ? = -42 + ?
3(x^2 - 4x + 4) + 5(y^2 + 6y + 9) = -42 + 3*4 + 5*9
3(x-2)^2 + 5(y+3)^2 = 15
(x-2)^2/5 + (y+3)^2/3 = 1
===============================
Cheers,
Stan H.